40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jan 14 03:42:31 2015, Jenne, Update, LSC, Thoughts on going forward with variable finesse 
    Reply  Thu Jan 15 04:41:01 2015, Jenne, Update, LSC, Thoughts on going forward with variable finesse 
       Reply  Thu Jan 15 14:28:14 2015, Jenne, Update, LSC, LSC model change idea SwitchableErrorSignalProposal_Jan2015.pdf
          Reply  Fri Jan 16 03:31:35 2015, Jenne, Update, LSC, LSC model change implemented LSCorig_15Jan2015.pngLSCpost_15Jan2015.png
             Reply  Fri Jan 16 18:46:15 2015, Koji, Update, LSC, LSC model change implemented 
                Reply  Fri Jan 16 20:01:32 2015, Jenne, Update, LSC, LSC model change implemented 
             Reply  Tue Jan 20 15:09:01 2015, Jenne, Update, LSC, LSC model change implemented 
             Reply  Sat Mar 7 18:53:48 2015, Jenne, Update, LSC, DoF selector matrices replaced with filter banks DoFselectors_7March2015.png
Message ID: 10910     Entry time: Fri Jan 16 03:31:35 2015     In reply to: 10904     Reply to this: 10914   10923   11113
Author: Jenne 
Type: Update 
Category: LSC 
Subject: LSC model change implemented 

Okay, it has taken me almost exactly 12 hours (with a dinner break), but I have implemented this change.

Everything was svn-ed before I did things, and then again afterward.

Here is the "before" screenshot of the LSC model:

And here is afterward:

If you look extra carefully, you will see that it matches my proposal from http://nodus.ligo.caltech.edu:8080/40m/10904 .  I have made the change for DARM, MICH, PRCL, SRCL and CARM.  I did not alter XARM, YARM or MC.  Also, the CESAR stuff was taken out of the CARM area, since this is now a more generalized version of the same thing.


I have also checked and modified all of the scripts that I could think of, as well as all of the ifoconfig burt .req and .snap files that I could think of.  I also ran the carm_cm_up.sh script once, and it seems to work fine.  All of the transition scripts that are listed below (which are the only ones used currently in the sequence) now use the new error signal blending scheme.

Scripts:

  • Lock_ALS_CARMandDARM.py
  • ALSfindIRresonance.py
  • ALSwatch.py
  • carm_cm_down.sh
  • carm_cm_up.sh
  • CheckPRMIlock.py
  • Transition_MICH_REFL33Q_to_ASDC.py
  • Transition_CARM_ALS_to_TransInvSqrt.py
  • Transition_DARM_ALS_to_DCtrans.py
  • UGFup.py
  • UGFdown.py

Burts (listed are the .req files, but I also checked the .snap files, hand-editing the matrix element numbers where needed if I wasn't in the right config to do a save):

  • C1configure_Yarm.req
  • C1configure_YarmALS.req
  • C1configure_Xarm.req
  • C1configure_XarmALS.req
  • C1configure_CARM.req
  • C1configure_DARM.req
  • C1configure_PRM_forCARMdarm.req
  • C1configure_PRM_SBres.req
  • C1configure_PRM_Carr.req
  • C1configure_PRY.req
  • C1configure_DRM.req
  • C1configure_SRM.req

I also modified the screens for the input matrix and for the normalization matrix.  I created a new screen for the final blending matrices (which are all 2x1's), and I also modified the LSC_OVERVIEW screen. 

The input matrix and normalization matrix screens have colored bars that tell you whether a row is in use or not.  If the background to the row is the blue of the whole screen, that row is not being used.

The LSC screen has new hand-modified Kissel Buttons.  I wanted to show the total PD error signal that is being used, regardless of what row (A or B) it is on at that time.  So, I have collapsed the rows so that DARM_A and DARM_B are overlapped, even though they are actually rows 1 and 2 of the matrix.  The PD should only show up green on the LSC screen if that row is in use (so, if you are prepping a row, but aren't using it yet, you won't see those elements in the matrix).  Anyhow, the point is that the LSC overview part of things shouldn't look any different than before.


Brain not working anymore now that it's ~4am, but I need to rethink and recheck to make sure that the PD whitening triggering is still okay and working.  Or maybe we can remove it, and just include that in the scripts, as Rana has been suggesting for ages.  Thoughts for tomorrow.

ELOG V3.1.3-